Benzene reacts with bromine to form bromobenzene, C6H5Br.
C6H6(L) + Br2(L) ---> C6H5Br(L) + HBr (g)

A) What is the maximum amount of C6H5Br that can be formed from the reaction of 7.50g of C6H6 with excess Br2?

B) A competing reaction is the formation of dibromobenzene, C6H4Br.
C6H6(l) + 2Br2(L) ----> C6H4Br2(L) + 2HBr(G)
If 1.25 g of C6H4Br2 was formed by the competing reaction, how much C6H6 was not coverted to C6H5Br?
1
Expert's answer
2019-03-18T02:20:55-0400

a) C6H6(l) + Br2(l) -> C6H5Br(l) + HBr(g)

n= m/M

n(C6H6) = 7.50 g / 78.11 g/mol = 0.096 mol

according to equation mole ratio n(C6H6):n(C6H5Br) = 1:1, then n(C6H5Br) = n(C6H6) = 0.096 mol

m(C6H6Br) = 0.096 mol * 157.01 g/mol =15.07 g


b) C6H6(l) + 2Br2(l) -> C6H4Br2(l) + 2HBr(g)

n= m/M

n(C6H4Br2) = 1.25 g/ 235.9 g/mol = 0.0053 mol

According to equation mole ratio n(C6H6):n(C6H4Br2) = 1:1, then n(C6H6) = n(C6H4Br2) = 0.0053 mol

m(C6H6) = 0.0053 mol * 78.11 g/mol = 0.412 g


Answer: a) 15.07 g; b) 0.412 g